# Streaming Operator Inference for Model Reduction of

### v1

@@ Appendix B Recursive Least-Squares Algorithms
The block-diagonal structure of $\mathbb{V}_{r}-\widetilde{\mathbb{V}}_{r}$ implies that its norm is the maximum norm of a diagonal block
@@
By assumption, $\|\mathbf{v}_{j}-\widetilde{\mathbf{v}}_{j}\|_{2}\leq\varepsilon_{v,j}$ for each $j$, so
$\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{F}\leq\Bigl(\sum_{j=1}^{r}\varepsilon_{v,j}^{2}\Bigr)^{1/2}=\tau_{v}.$
Since the Frobenius norm dominates the spectral norm, $\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\leq\tau_{v}$.
For the Kronecker product term, we add and subtract $\widetilde{\mathbf{V}}_{r}\otimes\mathbf{V}_{r}$ to obtain
@@
Applying the triangle inequality and the spectral norm property of Kronecker products [Lancaster and Farahat, 1972, Thm. 8], namely $\|\mathbf{A}\otimes\mathbf{B}\|_{2}=\|\mathbf{A}\|_{2}\|\mathbf{B}\|_{2}$, yields
$\|\mathbf{V}_{r}\otimes\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\otimes\widetilde{\mathbf{V}}_{r}\|_{2}\leq\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\|\mathbf{V}_{r}\|_{2}+\|\widetilde{\mathbf{V}}_{r}\|_{2}\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}=2\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\leq 2\tau_{v},$
where we used $\|\mathbf{V}_{r}\|_{2}=\|\widetilde{\mathbf{V}}_{r}\|_{2}=1$ since both matrices have orthonormal columns.
It remains to bound $\|\mathbb{D}\|_{2}$. Assuming bounded inputs $\|\mathbf{U}\|_{\infty}<\eta$ for some $\eta>0$, we have
@@
The Khatri–Rao product satisfies $(\mathbf{X}\odot\mathbf{X})^{\top}(\mathbf{X}\odot\mathbf{X})=(\mathbf{X}^{\top}\mathbf{X})\circ(\mathbf{X}^{\top}\mathbf{X})$, where $\circ$ denotes the Hadamard product. By [Horn and Johnson, 1994, Thm. 5.5.1], $\sigma_{1}((\mathbf{X}^{\top}\mathbf{X})\circ(\mathbf{X}^{\top}\mathbf{X}))\leq\sigma_{1}(\mathbf{X}^{\top}\mathbf{X})^{2}=\|\mathbf{X}\|_{2}^{4}$. Thus,
@@
where $\sigma_{1}=\|\mathbf{X}\|_{2}$. Substituting into (32) gives $\|\mathbf{E}_{D}\|_{2}\leq 2\beta_{1}^{\prime}\tau_{v}$. Note that $\beta_{1}^{\prime}=\beta_{1}/(2\sqrt{\min\{n,K\}})$ using the constant $\beta_{1}$ defined in (37), and hence, $\|\mathbf{E}_{D}\|_{2}\leq\beta_{1}\tau_{v}/\sqrt{\min\{n,K\}}$.
For the right-hand side perturbation, we have $\mathbf{E}_{R}=(\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r})^{\top}\dot{\mathbf{X}}$. Given the polynomial structure of the underlying dynamical system (1), the vector field is continuously differentiable and hence $L_{f}$-Lipschitz continuous for some $L_{f}>0$. This property can be used to bound $\|\dot{\mathbf{X}}\|_{2}$ in general settings. However, to obtain bounds that are directly comparable across paradigms and more readily interpretable, we assume the time-derivative data is computed via finite-difference approximations, i.e., $\dot{\mathbf{X}}=\mathbf{X}\boldsymbol{\Delta}_{K}$, where $\boldsymbol{\Delta}_{K}\in\mathbb{R}^{K\times K}$ is the finite-difference operator matrix defined in Section 3.1. This assumption is reasonable in practice, as time-derivative data is often unavailable and must be approximated from state snapshots. Under this assumption,
$\|\mathbf{E}_{R}\|_{F}=\|(\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r})^{\top}\mathbf{X}\boldsymbol{\Delta}_{K}\|_{F}\leq\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{F}\|\mathbf{X}\|_{2}\|\boldsymbol{\Delta}_{K}\|_{2}\leq\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v}.$
Reformulation paradigms (iSVD-LS/RLS)
For these paradigms, the data matrix perturbation takes the form
@@
Applying the triangle inequality, submultiplicativity, and the bounds $\|\mathbf{w}_{j}-\widetilde{\mathbf{w}}_{j}\|_{2}\leq\varepsilon_{w,j}$ and $|\sigma_{j}-\widetilde{\sigma}_{j}|\leq\varepsilon$ yields
$\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\|\boldsymbol{\Sigma}_{r}\|_{2}\|\mathbf{W}_{r}-\widetilde{\mathbf{W}}_{r}\|_{F}+\|\widetilde{\mathbf{W}}_{r}\|_{2}\|\boldsymbol{\Sigma}_{r}-\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\sigma_{1}\tau_{w}+\sqrt{r}\,\varepsilon=\phi_{w},$
where $\tau_{w}=(\sum_{j=1}^{r}\varepsilon_{w,j}^{2})^{1/2}$ and we used $\|\widetilde{\mathbf{W}}_{r}\|_{2}=1$. Since the Frobenius norm dominates the spectral norm, $\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{2}\leq\phi_{w}$.
For the Kronecker product term in (33), we proceed analogously to the projection case. Adding and subtracting $(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\otimes(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})$ and applying the triangle inequality and [Lancaster and Farahat, 1972, Thm. 8] gives
@@
Since $\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}\|_{2}=\sigma_{1}$ and $\|\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{2}=\max_{j}\widetilde{\sigma}_{j}\leq\sigma_{1}+\varepsilon$, we obtain
$\|(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})\otimes(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})-(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\otimes(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\|_{2}\leq(2\sigma_{1}+\varepsilon)\phi_{w}.$
Substituting into (33) along with $\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}=\beta_{2}/[\sqrt{r}(\sigma_{1}+\varepsilon)]$ from Equation 38 and expanding $\phi_{w}$ yields
$\|\mathbf{E}_{D}\|_{2}\leq\bigl[\phi_{w}^{2}+(2\sigma_{1}+\varepsilon)^{2}\phi_{w}^{2}\bigr]^{1/2}=\phi_{w}\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}=\sigma_{1}\beta_{2}\tau_{w}/[\sqrt{r}(\sigma_{1}+\varepsilon)]+\beta_{2}\varepsilon/(\sigma_{1}+\varepsilon).$
For the right-hand side perturbation, we have $\mathbf{R}=\boldsymbol{\Delta}_{K}^{\top}\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}$ and $\widetilde{\mathbf{R}}=\boldsymbol{\Delta}_{K}^{\top}\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}$, so
$\|\mathbf{E}_{R}\|_{F}=\|\boldsymbol{\Delta}_{K}^{\top}(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\|_{F}\leq\|\boldsymbol{\Delta}_{K}\|_{2}\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\|\boldsymbol{\Delta}_{K}\|_{2}\phi_{w}=\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{w}+\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}\varepsilon.\qed$
C.3 Proof of Theorem 3.3
To apply Lemma 3.1, we require bounds on $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\|_{2}$, $\|\widetilde{\mathbf{D}}^{\dagger}\|_{2}$, and $\|\widetilde{\mathbf{R}}\|_{F}$. We begin with a standard result relating the pseudoinverse norm to the minimum singular value.
Lemma C.2 (Norm of the pseudoinverse).
Proof.
Let $\mathbf{B}=\mathbf{A}^{\dagger}$. Since $\mathbf{A}\mathbf{B}$ is an orthogonal projector onto $\mathrm{range}(\mathbf{A})$, we have $\mathbf{A}\mathbf{B}\mathbf{x}\perp(\mathbf{I}-\mathbf{A}\mathbf{B})\mathbf{x}$ for any $\mathbf{x}\in\mathbb{R}^{n}$. Therefore,
$\|\mathbf{B}\|_{2}=\max_{\mathbf{x}\neq 0}\frac{\|\mathbf{B}\mathbf{x}\|_{2}}{\|\mathbf{x}\|_{2}}=\max_{\mathbf{x}\neq 0}\frac{\|\mathbf{B}\mathbf{x}\|_{2}}{\|\mathbf{A}\mathbf{B}\mathbf{x}+(\mathbf{I}-\mathbf{A}\mathbf{B})\mathbf{x}\|_{2}}=\max_{\mathbf{B}\mathbf{x}\neq 0}\frac{\|\mathbf{B}\mathbf{x}\|_{2}}{\|\mathbf{A}\mathbf{B}\mathbf{x}+(\mathbf{I}-\mathbf{A}\mathbf{B})\mathbf{x}\|_{2}}\leq\max_{\mathbf{B}\mathbf{x}\neq 0}\frac{\|\mathbf{B}\mathbf{x}\|_{2}}{\|\mathbf{A}\mathbf{B}\mathbf{x}\|_{2}},$
where the inequality follows from orthogonality. Substituting $\mathbf{y}=\mathbf{B}\mathbf{x}$ and noting that $\mathbf{y}$ ranges over $\mathrm{range}(\mathbf{B})=\mathrm{range}(\mathbf{A}^{\top})$ gives
$\|\mathbf{B}\|_{2}\leq\max_{\mathbf{y}\in\mathrm{range}(\mathbf{A}^{\top}),\mathbf{y}\neq 0}\frac{\|\mathbf{y}\|_{2}}{\|\mathbf{A}\mathbf{y}\|_{2}}=\Bigl(\min_{\mathbf{y}\in\mathrm{range}(\mathbf{A}^{\top}),\mathbf{y}\neq 0}\frac{\|\mathbf{A}\mathbf{y}\|_{2}}{\|\mathbf{y}\|_{2}}\Bigr)^{-1}=\sigma_{\min}(\mathbf{A})^{-1}.$
Equality is attained by taking $\mathbf{x}$ in the direction of the right singular vector corresponding to $\sigma_{\min}(\mathbf{A})$. ∎
Bounds on $\|\widetilde{\mathbf{R}}\|_{F}$
For the projection-based paradigms, $\widetilde{\mathbf{R}}=\widetilde{\mathbf{V}}_{r}^{\top}\dot{\mathbf{X}}$. Under the finite-difference assumption $\dot{\mathbf{X}}=\mathbf{X}\boldsymbol{\Delta}_{K}$,
For the reformulation paradigms, $\widetilde{\mathbf{R}}=\boldsymbol{\Delta}_{K}^{\top}\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}$, so
@@
$\|\widetilde{\mathbf{R}}\|_{F}\leq\|\boldsymbol{\Delta}_{K}\|_{2}\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F},$ (35)
@@
since $\widetilde{\mathbf{W}}_{r}$ has orthonormal columns. To bound $\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}$, let $\sigma=[\sigma_{1},\ldots,\sigma_{r}]^{\top}$ and $\delta=[\delta_{1},\ldots,\delta_{r}]^{\top}$ where $\delta_{j}=\widetilde{\sigma}_{j}-\sigma_{j}$ satisfies $|\delta_{j}|\leq\varepsilon$. Then using the inner product operator $\langle\cdot,\cdot\rangle$ we have
$\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}^{2}=\sum_{j=1}^{r}(\sigma_{j}+\delta_{j})^{2}=\|\boldsymbol{\Sigma}_{r}\|_{F}^{2}+2\langle\sigma,\delta\rangle+\|\delta\|_{2}^{2}.$
@@
The Cauchy-Schwarz inequality gives $|\langle\sigma,\delta\rangle|\leq\|\sigma\|_{2}\|\delta\|_{2}\leq\|\boldsymbol{\Sigma}_{r}\|_{F}\cdot\sqrt{r}\varepsilon$, and $\|\delta\|_{2}\leq\sqrt{r}\varepsilon$. Therefore,
@@
Bounds on $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\|_{2}$ and $\|\widetilde{\mathbf{D}}^{\dagger}\|_{2}$
In OpInf, the data matrix $\mathbf{D}$ is commonly ill-conditioned due to linear dependence among columns arising from small time-step sizes and/or the quadratic nonlinearity Peherstorfer and Willcox [2016], Geelen and Willcox [2022]. When the state data exhibits rapid spectral decay, the quadratic terms amplify this effect by squaring the singular values, further reducing $\sigma_{\min}(\mathbf{D})$. Consequently, it is often necessary to choose the regularization parameter such that $\sqrt{\gamma_{\min}}=\min_{j}\sqrt{\gamma_{j}}\gg\sigma_{\min}(\mathbf{D})$. With a similar reasoning on $\widetilde{\mathbf{D}}$, we suppose the perturbed minimum singular value satisfies $\sqrt{\gamma_{\min}}\gg\sigma_{\min}(\mathbf{D}+\mathbf{E}_{D})$.
Under these assumptions, Lemma C.2 gives
$\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\|_{2}=\sigma_{\min}(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu)^{-1}=\bigl[\sigma_{\min}(\mathbf{D})^{2}+\gamma_{\min}\bigr]^{-1/2}\approx\gamma_{\min}^{-1/2},$
and we similarly have
$\|\widetilde{\mathbf{D}}^{\dagger}\|_{2}=\bigl[\sigma_{\min}(\mathbf{D}+\mathbf{E}_{D})^{2}+\gamma_{\min}\bigr]^{-1/2}\approx\gamma_{\min}^{-1/2}.$
@@
Substituting the bounds from Lemma 3.2 and the estimates above into Lemma 3.1, we obtain for the projection paradigms (iSVD-Project-LS/RLS):
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}\leq\frac{\alpha}{\gamma_{\min}}\cdot 2\beta_{1}^{\prime}\tau_{v}\cdot\sigma_{1}\sqrt{\min\{n,K\}}\|\boldsymbol{\Delta}_{K}\|_{2}+\frac{1}{\sqrt{\gamma_{\min}}}\cdot\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v}=\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{v}\Bigl(1+\frac{\alpha\beta_{1}}{\sqrt{\gamma_{\min}}}\Bigr),$ (37)
where $\beta_{1}=2\beta_{1}^{\prime}\sqrt{\min\{n,K\}}$.
For the reformulation paradigms (iSVD-LS/RLS), substituting the bounds yields
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}$ $\leq\frac{\alpha}{\gamma_{\min}}\cdot\phi_{w}\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}\cdot\|\boldsymbol{\Delta}_{K}\|_{2}\sqrt{r}(\sigma_{1}+\varepsilon)+\frac{1}{\sqrt{\gamma_{\min}}}\cdot\|\boldsymbol{\Delta}_{K}\|_{2}\phi_{w}.$
Expanding $\phi_{w}=\sigma_{1}\tau_{w}+\sqrt{r}\varepsilon$ and collecting terms gives
where $\beta_{2}=\sqrt{r}(\sigma_{1}+\varepsilon)\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}$ and both $\beta_{1},\beta_{2}$ have order $O(\sigma_{1}^{2})$ if $\sigma_{1}$ dominates. This completes the proof.

### v2

@@ C.2 Operator Error Bounds for Each Paradigm
We first bound $\|\mathbb{V}_{r}-\widetilde{\mathbb{V}}_{r}\|_{2}$. The block-diagonal structure implies [Tropp, 2016, p. 17]
@@
Since the Frobenius norm dominates the spectral norm and $\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{F}=\tau_{v}$, we have $\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\leq\tau_{v}$. For the Kronecker product term, we add and subtract $\widetilde{\mathbf{V}}_{r}\otimes\mathbf{V}_{r}$ to obtain
@@
The triangle inequality and the spectral norm property $\|\mathbf{A}\otimes\mathbf{B}\|_{2}=\|\mathbf{A}\|_{2}\|\mathbf{B}\|_{2}$ [Lancaster and Farahat, 1972, Thm. 8] yield
$\|\mathbf{V}_{r}\otimes\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\otimes\widetilde{\mathbf{V}}_{r}\|_{2}$ $\leq\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\|\mathbf{V}_{r}\|_{2}+\|\widetilde{\mathbf{V}}_{r}\|_{2}\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}$
$=2\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{2}\leq 2\tau_{v},$
where we used $\|\mathbf{V}_{r}\|_{2}=\|\widetilde{\mathbf{V}}_{r}\|_{2}=1$ since both matrices have orthonormal columns. Thus, $\|\mathbb{V}_{r}-\widetilde{\mathbb{V}}_{r}\|_{2}\leq 2\tau_{v}$.
Next, we bound $\|\mathbb{D}\|_{2}$. With bounded inputs $\|\mathbf{U}\|_{2}=\eta$ for some $\eta>0$, we have
@@
For the Khatri-Rao product term, we use the identity $(\mathbf{X}\odot\mathbf{X})^{\top}(\mathbf{X}\odot\mathbf{X})=(\mathbf{X}^{\top}\mathbf{X})\circ(\mathbf{X}^{\top}\mathbf{X})$, where $\circ$ denotes the Hadamard product. By [Horn and Johnson, 1994, Thm. 5.5.1], $\sigma_{1}((\mathbf{X}^{\top}\mathbf{X})\circ(\mathbf{X}^{\top}\mathbf{X}))\leq\sigma_{1}(\mathbf{X}^{\top}\mathbf{X})^{2}=\|\mathbf{X}\|_{2}^{4}$. Therefore,
@@
where $\sigma_{1}=\|\mathbf{X}\|_{2}$. Substituting into (31) gives
$\|\mathbf{E}_{D}\|_{2}\leq 2\beta_{1}^{\prime}\tau_{v}=\frac{\beta_{1}}{\sqrt{\min\{n,K\}}}\tau_{v},$ (32)
where $\beta_{1}=2\beta_{1}^{\prime}\sqrt{\min\{n,K\}}$.
Bounding the right-hand side perturbation $\mathbf{E}_{R}$
For the projection paradigms, $\mathbf{E}_{R}=(\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r})^{\top}\dot{\mathbf{X}}$. Using the finite-difference assumption $\dot{\mathbf{X}}=\mathbf{X}\boldsymbol{\Delta}_{K}$ and submultiplicativity,
$\|\mathbf{E}_{R}\|_{F}=\|(\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r})^{\top}\mathbf{X}\boldsymbol{\Delta}_{K}\|_{F}\leq\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{F}\|\mathbf{X}\|_{2}\|\boldsymbol{\Delta}_{K}\|_{2}=\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v}.$ (33)
Bounding the pseudoinverse norms
In OpInf, the data matrix $\mathbf{D}$ is commonly ill-conditioned due to linear dependence among columns arising from small time steps and the quadratic nonlinearity Peherstorfer and Willcox [2016], Geelen and Willcox [2022]. When the state data exhibits rapid spectral decay, quadratic terms amplify this effect by squaring the singular values, further reducing $\sigma_{\min}(\mathbf{D})$. The regularization parameter is typically chosen to satisfy $\sqrt{\gamma_{\min}}=\min_{j}\sqrt{\gamma_{j}}\gg\sigma_{\min}(\mathbf{D})$. Similar reasoning applies to $\widetilde{\mathbf{D}}$, giving $\sqrt{\gamma_{\min}}\gg\sigma_{\min}(\mathbf{D}+\mathbf{E}_{D})$.
Hence, Lemma C.2 yields
$\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\|_{2}=\sigma_{\min}(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu)^{-1}=\bigl[\sigma_{\min}(\mathbf{D})^{2}+\gamma_{\min}\bigr]^{-1/2}\approx\gamma_{\min}^{-1/2},$ (34)
and similarly,
$\|\widetilde{\mathbf{D}}^{\dagger}\|_{2}=\bigl[\sigma_{\min}(\mathbf{D}+\mathbf{E}_{D})^{2}+\gamma_{\min}\bigr]^{-1/2}\approx\gamma_{\min}^{-1/2}.$ (35)
For $\|\widetilde{\mathbf{R}}\|_{F}$, we use $\widetilde{\mathbf{R}}=\widetilde{\mathbf{V}}_{r}^{\top}\dot{\mathbf{X}}=\widetilde{\mathbf{V}}_{r}^{\top}\mathbf{X}\boldsymbol{\Delta}_{K}$ to obtain
Completing the proof
Substituting the bounds from (32), (33), (34), (35), and (36) into Lemma 3.1 yields
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}$ $\leq\frac{\alpha}{\gamma_{\min}}\cdot\frac{\beta_{1}}{\sqrt{\min\{n,K\}}}\tau_{v}\cdot\sigma_{1}\sqrt{\min\{n,K\}}\|\boldsymbol{\Delta}_{K}\|_{2}+\frac{1}{\sqrt{\gamma_{\min}}}\cdot\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v}=\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{v}\Bigl(1+\frac{\alpha\beta_{1}}{\sqrt{\gamma_{\min}}}\Bigr),$
which completes the proof. ∎
C.2.2 Proof of Theorem 3.3
Proof.
We follow a similar strategy as for Theorem 3.2, deriving bounds on $\mathbf{E}_{D}$ and $\mathbf{E}_{R}$ specific to the reformulation paradigms.
Bounding the left-hand side perturbation $\mathbf{E}_{D}$
For the reformulation paradigms, the data matrix perturbation has the block structure
@@
Applying the triangle inequality, submultiplicativity, the relation $\|\mathbf{W}_{r}-\widetilde{\mathbf{W}}_{r}\|_{F}=\tau_{w}$, and Lemma C.3 yields
$\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\sigma_{1}\|\mathbf{W}_{r}-\widetilde{\mathbf{W}}_{r}\|_{F}+\|\widetilde{\mathbf{W}}_{r}\|_{2}\|\boldsymbol{\Sigma}_{r}-\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\sigma_{1}\tau_{w}+\sqrt{r}\,\varepsilon=:\phi_{w},$ (38)
where we used $\|\widetilde{\mathbf{W}}_{r}\|_{2}=1$ and $\|\boldsymbol{\Sigma}_{r}-\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}\leq\sqrt{r}\varepsilon$. Since the Frobenius norm dominates the spectral norm, $\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{2}\leq\phi_{w}$.
For the Kronecker product term in (37), we add and subtract $(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\otimes(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})$ and apply the triangle inequality and [Lancaster and Farahat, 1972, Thm. 8]:
@@
Since $\|\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}\|_{2}=\sigma_{1}$ and $\|\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}\|_{2}=\max_{j}\widetilde{\sigma}_{j}\leq\sigma_{1}+\varepsilon$ by Lemma C.3, we obtain
$\|(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})\otimes(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r})-(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\otimes(\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\|_{2}\leq(2\sigma_{1}+\varepsilon)\phi_{w}.$ (39)
Substituting (38) and (39) into (37) and taking square roots gives
$\|\mathbf{E}_{D}\|_{2}\leq\phi_{w}\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}.$
Expanding $\phi_{w}=\sigma_{1}\tau_{w}+\sqrt{r}\varepsilon$ and defining $\beta_{2}=\sqrt{r}(\sigma_{1}+\varepsilon)\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}$ yields
$\|\mathbf{E}_{D}\|_{2}=\frac{\sigma_{1}\beta_{2}}{\sqrt{r}(\sigma_{1}+\varepsilon)}\tau_{w}+\frac{\beta_{2}}{\sigma_{1}+\varepsilon}\varepsilon.$ (40)
Bounding the right-hand side perturbation $\mathbf{E}_{R}$
For reformulation paradigms, $\mathbf{R}=\boldsymbol{\Delta}_{K}^{\top}\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}$ and $\widetilde{\mathbf{R}}=\boldsymbol{\Delta}_{K}^{\top}\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}$, so
$\|\mathbf{E}_{R}\|_{F}=\|\boldsymbol{\Delta}_{K}^{\top}(\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}-\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r})\|_{F}\leq\|\boldsymbol{\Delta}_{K}\|_{2}\phi_{w}=\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{w}+\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}\varepsilon.$ (41)
Bounding $\|\widetilde{\mathbf{R}}\|_{F}$
@@
Since $\widetilde{\mathbf{R}}=\boldsymbol{\Delta}_{K}^{\top}\widetilde{\mathbf{W}}_{r}\widetilde{\boldsymbol{\Sigma}}_{r}$ and $\widetilde{\mathbf{W}}_{r}$ has orthonormal columns,
$\|\widetilde{\mathbf{R}}\|_{F}\leq\|\boldsymbol{\Delta}_{K}\|_{2}\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}.$ (42)
@@
To bound $\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}$, let $\boldsymbol{\sigma}=[\sigma_{1},\ldots,\sigma_{r}]^{\top}$ and $\boldsymbol{\delta}=[\delta_{1},\ldots,\delta_{r}]^{\top}$ where $\delta_{j}=\widetilde{\sigma}_{j}-\sigma_{j}$ satisfies $|\delta_{j}|\leq\varepsilon$ by Lemma C.3. Then
$\|\widetilde{\boldsymbol{\Sigma}}_{r}\|_{F}^{2}=\sum_{j=1}^{r}(\sigma_{j}+\delta_{j})^{2}=\|\boldsymbol{\Sigma}_{r}\|_{F}^{2}+2\boldsymbol{\sigma}^{\top}\boldsymbol{\delta}+\|\boldsymbol{\delta}\|_{2}^{2}.$
@@
The Cauchy-Schwarz inequality gives $|\boldsymbol{\sigma}^{\top}\boldsymbol{\delta}|\leq\|\boldsymbol{\sigma}\|_{2}\|\boldsymbol{\delta}\|_{2}\leq\|\boldsymbol{\Sigma}_{r}\|_{F}\cdot\sqrt{r}\varepsilon$, and $\|\boldsymbol{\delta}\|_{2}\leq\sqrt{r}\varepsilon$. Therefore,
@@
The bounds for $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\|_{2}$ and $\|\widetilde{\mathbf{D}}^{\dagger}\|_{2}$ follow from (34) and (35) as in the projection case. Substituting the expressions from (40), (41), (34), (35), and (43) into Lemma 3.1 gives
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}$ $\leq\frac{\alpha}{\gamma_{\min}}\cdot\Biggl(\frac{\sigma_{1}\beta_{2}}{\sqrt{r}(\sigma_{1}+\varepsilon)}\tau_{w}+\frac{\beta_{2}}{\sigma_{1}+\varepsilon}\varepsilon\Biggr)\cdot\|\boldsymbol{\Delta}_{K}\|_{2}\sqrt{r}(\sigma_{1}+\varepsilon)+\frac{1}{\sqrt{\gamma_{\min}}}\cdot\|\boldsymbol{\Delta}_{K}\|_{2}(\sigma_{1}\tau_{w}+\sqrt{r}\varepsilon).$
Collecting terms and simplifying yields
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}\leq\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{w}\Bigl(1+\frac{\alpha\beta_{2}}{\sqrt{\gamma_{\min}}}\Bigr)+\frac{(\alpha\beta_{2}+\sqrt{\gamma_{\min}})\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}}{\gamma_{\min}}\varepsilon,$
which completes the proof. ∎